OpenStudy (anonymous):

if y=cos(2x), determine d^80y/dx^80

OpenStudy (tkhunny):

It's cyclical-ish. 1) (d/dx) cos(2x) = -2sin(2x) 2) (d/dx) -2sin(2x) = -4cos(2x) 3) (d/dx) -4cos(2x) = 8sin(2x) 4) (d/dx) 8sin(2x) = 16cos(2x) No the thinking about it. \[2^{4}cos(2x)\] After 8 \[2^{8}cos(2x)\] After 16 \[2^{16}cos(2x)\] After 32 \[2^{32}cos(2x)\] Do you see it?

OpenStudy (anonymous):

Yes, so it's \[2^{80}\sin(2x)\] Thank you~ :D

OpenStudy (tkhunny):

Why did you switch it to sine? Any multiple of 4 will be a positive cosine.

OpenStudy (anonymous):

ohh... when would it switch to sine?

OpenStudy (tkhunny):

That's why I wrote out the first 4. 1 more than a multiple of 4 leads to negative sine 2 more than a multiple of 4 leads to negative cosine 3 more than a multiple of 4 leads to positive sine Multiple of 4 leads to positive cosine

OpenStudy (anonymous):

Thank you again, I'll write it down for future reference. If it goes onto 5, would that be back to negative sine?

OpenStudy (tkhunny):

Again, that's why I wrote it all that way. I did not stick to a specific case. I wrote it for ALL possibilities. Read it more carefully and you will see it.